OpenStudy (anonymous):

is the sequence arithmetic? if so, identify the common difference. 10, 14, 18, 22,...

jimthompson5910 (jim_thompson5910):

Take each term and subtract off the previous term 14 - 10 = ??? 18 - 14 = ??? 22 - 18 = ???

jimthompson5910 (jim_thompson5910):

If each difference is the same number, then you have an arithmetic sequence and the common difference is equal to those differences

OpenStudy (ksaimouli):

(a=10 and r=4) ya u need to subtract